A small particle moves horizontally at 3.0 x10^6 m/s and entered a downward electric field produced by two horizontal plates, 2.0 mm apart, connected to a 40-V battery. If the charge-to-mass ratio of the particle is 1.0 x10^8 C/kg, what is the instantaneous acceleration?
1
Expert's answer
2018-06-04T08:27:08-0400
F=ma
F=qE=qU/d.
Thus,
ma=qU/d
a=q/m U/d=1.0∙〖10〗^8 40/0.002=2.0∙〖10〗^12 m/s^2